# The building blocks of software work explain coding careers and language popularity

### v1

@@ C.1 Example job ad
The analyses in Fig. 2 of the main text, we assign SO tasks to job advertisement. To do so, we first feed the job ad’s text to a large language model (ChatGPT-3.5) to extract strings that describe all task requirements from the listed job description. Next, we use sBert[Reimers and Gurevych, 2019] to embed each of the job ad’s task requirements in an 384 dimensional space. We then do the same for SO task labels we generated in Sec. B. To increase the accuracy of these embeddings we prompt ChatGPT-4.0 to produce somewhat longer labels of up to 30 words. Finally for each task requirement, we pick the closest SO task in terms of the cosine similarity between the two embeddings. We keep only those task requirements for which the matched SO task’s cosine similarity exceeds 0.3.
Below, we illustrate this using a concrete example of a job advertisement. The job advertisement contains the following text:
@@
The task requirements the LLM extracts, together with the SO tasks we match to them, are listed in Table C.1. The procedure leads to plausible matches. The first job requirement, AWS (Amazon Web Services, a cloud computing service), is clearly closely related to the matched SO task, which contains several tags related to scheduling tasks, such as job-scheduling or quartz-scheduler and several amazon-specific tags such as aws-application-load-balancer and amazon-athena. Also for the second job requirement, the identified match is plausible. Ansible and Terraform are two platforms for CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines, which help automate the building, testing, and deployment of code. In this context, Maven is primarily used in Java and Kubernetes focuses on automating the deployment of containerized applications. For the third and fourth job requirements, the similarity to their matched SO tasks falls below the threshold value of 0.3. Yet, the job requirements and matched SO tasks are still related, albeit somewhat more weakly. In the third job requirement, metric aggregation systems are often used to monitor CPU or network loads, which support scalable real-time data processing. Similary, the PCI (Payment Card Industry compliance mentioned in the fourth job requirement covers security standards required when handling debit or credit card information. This typically involves the implementation of secure user authentication and authorization processes of the matched SO task.
Table C.1: Tasks requirements in job ads and their matching SO tasks.
(1) task requirements (2) matched SO task similarity (cosine)
@@
Column (1): required tasks as extracted by the LLM from the job ad. Column (2): the label of the matched SO tasks. Column (3): cosine similarity between the embedding vectors of Columns (1) and (2).
C.2 Salary versus task value
To assess how well SO tasks predict salaries in job ads, we focus on the subset of job advertisements that also provide salary information. For these jobs, we calculate the total number of SO tasks they list, as well as the average value of these tasks. Next, we estimate the following regression model:
@@ C.2 Salary versus task value
where $w_{j}$ is the salary or midpoint of the salary range posted for job $j$, $\bar{v}_{j}$ is the average of the logarithm of the task values across all SO tasks job $j$ requires, $\bar{r}_{j}$ the average of the task density around these SO tasks (i.e., the coherence of the tasks the job requires) and $n_{j}$ the number of SO tasks job $j$ requires. $\tau_{y(j)}$ represents year fixed effects to control for general trends in the labor market. Finally, $\varepsilon_{j}$ is an error term.
Outcomes are reported in the table of Table C.2. The estimated effect effect of task value on posted salaries is around 0.9 across all models, suggesting that a $10\%$ increase in task value is associated with a $9\%$ higher salary.
@@
Robust standard errors in parentheses. p-value: ***: 0.01, **: 0.05, *: 0.1. The dependent variable is the value or midpoint of the salary range posted in a job ad, task value is the average of the logarithm of the task values across all SO tasks the job requires, task coherence is the average of the task density round all SO tasks the job requires, # tasks is the logarithm of the number of SO tasks the job requires.
C.3 Alternative job task matches and thresholds
@@
We test three alternative methods to match task requirements in job ads to SO tasks. To do so, let $v_{t}^{j}$ denote the embedding vector of task requirement $t$ of job $j$. Our first method, used in the main text, calculates the cosine similarity between $v_{t}^{j}$ and the embedding vector of each of the SO task’s (long) labels, and then selects the SO task that yields the highest cosine similarity. Our second method uses the similarity of $v_{t}^{j}$ to the embedding of the most prominent tag (in terms of its usage on SO) in each SO task. Our third method calculates the similarity of $v_{t}^{j}$ to the average embeddings of the tags defining the SO task. Finally, our fourth method uses the similarity of $v_{t}^{j}$ to the embedding of the closest tag in each SO task. Next, we repeat the analysis depicted in Fig. 2 with each of these methods. Figure C.1 presents the results.
@@
Figure C.1: Alternative methods to match task requirements in job ads to SO tasks. a,c,e,g. Prediction of task requirements. $20\%$ of task vector elements are masked and grouped into 10 equally sized bins based on the average relatedness of the masked task to the (unmasked) SO task requirements of a job. Plot displays the estimated probability, with its $95\%$ confidence interval, that tasks in a bin are required in the job. b,d,f,h. Predictions of wage offers. Jobs are grouped into equally sized bins based on the average value of the tasks they require. Vertical axis shows the estimated mean salary posted in the job ads, with their $95\%$ confidence intervals. a,b. Task requirements are matched to SO tasks using the embedding vector of the SO task’s label, namely the results shown in Fig 2. c,d. Task requirements are matched to SO tasks using the embedding vector of the SO task’s main (most frequent) tag. e,f. Task requirements are matched to SO tasks using the average embedding vector across all of the SO task’s tags. g,h. Task requirements are matched to SO tasks using the closest embedding vector across all tags of SO tasks.
@@
In the main text, we only keep task requirements that can be matched with SO tasks at a cosine similarity of $0.3$ or above. To explore the significance of this threshold, we repeat our analyses with two another thresholds: $0.2$ and $0.4$. The results are shown in Fig. C.2. Varying the threshold around the $0.3$ value of the main text does not meaningfully alter the outcomes of this analysis: density around masked tasks is strongly predictive of jobs demanding these tasks and salaries posted in job ads rise with the average values of the SO tasks we identify for them.
@@
Figure C.2: Results of task and salary prediction at varying matching thresholds. a,c,e. Probability that a masked SO task is listed among a job’s task requirements. Masked tasks represent $40\%$ of task vector elements and grouped into 10 equally sized bins based on the average relatedness of (density around) the masked task to the (unmasked) task requirements of a job. Plot displays estimated probabilities in each bin. b,d,f Predictions of salaries. Jobs are grouped into equally sized bins based on the average value of their required matched SO tasks. Vertical axis shows the estimated means in each bin of the advertised wage offers. a,b. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.3$, namely the results shown in Fig. 2. c,d. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.2$. e,f. Task requirements are matched to SO tasks with a minimum cosine similarity of $0.4$. Vertical bars represent 95% confidence intervals.
@@ D.1 Alternative specifications for voting regression
In Fig. 3c of the main text, we study how the popularity of an answer relates to the task experience of the user that provided it. The figure shows that answers by users with more task experiences tend to receive more upvotes and are more likely to be the top answer to a question. Table D.1 provides the full set of regression results related to this analysis.
@@
Regression model of eq. (8). The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. task experience: logarithm of task experience, the total number of answers the users provided to questions related to the focal task in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain either year or task-year fixed effects.
Because for questions that receive only a single answer, this answer is automatically the top answer, we repeat the analysis in a sample that excludes any question that receives only one answer. Table D.2 shows that this leads to larger observed effects of task experience.
@@ D.2 Causal effects: Instrumental variable estimation
In the main text, we raise the issue of sample selection bias in the regression analysis of the number of upvotes an answer receives on SO. There are two concerns. First, users can observe all existing answers to a question and, arguably, users will take these answers into consideration before deciding whether or not to provide an answer themselves. As a consequence, the answers we observe on SO are no random sample of user-answer combinations. For instance, we expect that users who believe they have little to add to the existing answers to a given question will tend to refrain from posting their own answers. This bias is addressed by only focusing on the first answer (in a temporal sense) that each question received.
@@
Second, even the first answer to a question will not be randomly sampled from a population of users and the answers they would have provided had they chosen to do so. Instead, answers will be provided by users who believe they have useful insights to the problem at hand. To address the sample selection bias this causes, we will exploit the fact that, due to timezone differences, most users are active on the SO platform at somewhat predictable times. Therefore, we expect that who will provide the first answer to a question to some extent depends on the time of day the question was posted. This introduces some exogenous variation in the likelihood that a given user provides an answer to a given question. Moreover, as far as task experience is not fully equally distributed across timezones, this offers exogenous variation in the task experience associated with the first answer to a question.
@@
We proceed as follows. First, we label each minute of a day, henceforth day-minute, from 0 to 1399 in a manner that is independent of timezones. Next, for each user, we calculate the average minute of the day they post answers on SO. This offers a coarse indication of when a user is active on the platform. Finally, we calculate for each task and each day-minute the amount of task experience that is expected to be available on SO. To do so, we sum the task experience of all users, weighted by the gap between the day-minute of a question and the average day-minute of each user. We call this variable, $M_{\theta,m}$, the minute-task count of task $\theta$ at day-minute $m$. Next, we use this variable as an instrument for the actual task experience of whichever user provides the first answer to the question at hand.
@@
Note that there are some potential concerns about this instrument. First, if there are geographic components to task experience of users and if these geographic components directly influence the quality of an answer, the instrument may be invalid. For instance, the timezone that includes Silicon Valley may have a disproportionate number of highly skilled programmers, such that any questions posed when most people in this timezone are active may attract high quality first answers. To correct for this, we always include day-minute fixed effects and in some specifications also add task-year-combination fixed effects.
@@
Second, the instrument assumes that questions are answered within day from the time they are posted on SO. In fact, we would expect that our instrument does not work for questions whose first answer does not arrive within 24 hours. We exploit this in placebo tests that focus on questions whose first answer is provided at least 24 hours after the posting of the question itself.
@@
where $u(a)$ is the user who posts answer $a$, $X_{\theta,u(a)}$ user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers that will eventually be provided to question $q(a)$, $Q_{a}$ the set of these answers, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term. Compared to eq. (8) in the main text, we add day-minute fixed effects, $\mu_{(q_{a}})$, that control for any timezone specific confounders.
As before, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is the top answer to its question or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. If an answer is associated with multiple tasks, we replicate the observation accordingly.
The results is presented in Table D.3. The effect of task experience on answer popularity is significant and positive across all model specifications. Moreover, the estimated causal effects arrived at by instrumental variables estimation are substantially larger than the OLS estimates reported in the main text. This is in line with a downward sample-selection bias in the OLS results.
@@
Clustered (group_task_qminute) standard-errors in parentheses
@@
Results of eq. (15) estimated with 2SLS. The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. task experience: logarithm of task experience, the total number of answers the user provided to questions related to the focal task in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain day-minute fixed effects (QA-abs-minute) for the timing of the question and year or task-year fixed effects. To overcome sample-selection biases, we instrument task experience with $M_{\theta,m}$, the estimated amount of available task experience at the moment the question was posted, calculated as the weighted sum of task experience across all users, where weights reflect the temporal distance between the day-minute of a user and of the question’s posting.
Next, we test this strategy using our placebo tests. To do so, we split the sample of questions into questions that receive a first answer within the first 24 hours after they were posted and those that receive their first answer later than that. Results are shown in Table D.4. For answers provided within 24 hours, task experience has a strong and statistically significant positive effect on answer popularity, regardless of whether we look at the probability of being the top answer or the number of upvotes an answer receives. In contrast, for the placebo test, comprising of answers provided more than 24 hours after the question was posted, the effect of task experience becomes insignificant in three out of four models. In the model without task fixed effects, the effect on the number of votes an answer receives turns significantly negative. This suggests that the task fixed effects control for some relevant confounders. Therefore, our preferred estimates are derived from models that include these fixed effects. These models suggest a large and positive effect of task experience on the popularity of answers, corroborating our analysis in the main text.
@@
Clustered (group_task_qminute) standard-errors in parentheses
@@
D.3 Regression of entry on task value
To study how task values effect users’ adoption on new tasks, we run the regression described in eq. (9) of the main text. We compare the association of two variables, task value and task density, with the likelihood that users adopt a certain task. We repeat this analysis once using all answers to construct experience and task entry variables, and once using only answers to Python related questions. The results are shown in Table D.5.
Table D.5: Task entry regressions
Dependent Variable: entry (binary, all programming languages) entry (binary, Python)

### v2

@@ C.1 Example job ad
The analyses in Fig. 2 of the main text, we assign skills to job advertisement. To do so, we first feed the job ad’s text to a large language model (ChatGPT-3.5) to extract strings that describe all work activities (“tasks”) and salary information from the listed job description, with the following prompt:
Here is a job ad:{text}. Could you extract in a json format the following information with required json key names:
1. company name, with key name ‘company_name’.
2. bool of whether the ad is for more than one job, with key name ’multi_jobs’.
3. With key name ’jobs’, build a list containing the information of each jobs: for each job, build a key-values pair with the job name as key name and value containing the following items: bool of whether this ad contain information of yearly salary values, with key name ’salary_index’; location, with key name ’location’; salary, with key name ’salary’ and containing the following items: if the salary is a range, please give an average value with the key name ’salary_average’, the minimum value with the key name ’salary_min’, the maximum value with the key name ’salary_max’; if the salary is not a range, please give the value with the key name ’salary_average’; currency of salary, with key name ’currency’; bool of whether salary includes equity, with key name ’include_equity’; a list of skills required by the job, with key name ’skills’.
Please only use the content in the advertisement and do not use other outside information.
Next, we use sBert[Reimers and Gurevych, 2019] to embed each of the listed work activities in a 384 dimensional space. We then do the same for the skill labels we generated in Sec. B. To increase the accuracy of these embeddings we prompt ChatGPT-4.0 to produce somewhat longer labels of up to 30 words. Finally for each work activity, we pick the closest skill in our taxonomy in terms of the cosine similarity between the two embeddings. We keep only those work activities for which the matched skill’s cosine similarity exceeds 0.3.
Below, we illustrate this using an example of a concrete job advertisement. The job advertisement contains the following text:
@@
The job requirements the LLM extracts, together with the skills we match to them, are listed in Table C.1. The procedure leads to plausible matches. The first job requirement, AWS (Amazon Web Services, a cloud computing service), is clearly closely related to the matched skill, which contains several tags related to scheduling tasks, such as job-scheduling or quartz-scheduler and several amazon-specific tags such as aws-application-load-balancer and amazon-athena. Also for the second job requirement, the identified match is plausible. Ansible and Terraform are two platforms for CI/CD (Continuous Integration and Continuous Delivery/Deployment) pipelines, which help automate the building, testing, and deployment of code. In this context, Maven, which is primarily used in Java and Kubernetes, focuses on automating the deployment of containerized applications. For the third and fourth job requirements, the similarity to their matched skills falls below the threshold value of 0.3. Yet, the job requirements and matched skills are still related, albeit somewhat more weakly. In the third job requirement, metric aggregation systems are often used to monitor CPU or network loads, which support scalable real-time data processing. Similarly, the PCI (Payment Card Industry) compliance mentioned in the fourth job requirement covers security standards required when handling debit or credit card information. This typically involves the expertise in implementing secure user authentication and authorization processes related to the matched skill.
Table C.1: Matching job requirements to skills.
(1) job requirements (2) matched skill similarity (cosine)
@@
Column (1): job requirements as extracted by the LLM from the job ad. Column (2): the label of the matching skills. Column (3): cosine similarity between the embedding vectors of Columns (1) and (2).
C.2 Salary versus skill value
To assess how well our skill taxonomy can be used to predict salaries in job ads, we focus on the subset of job advertisements that also provide salary information. For these jobs, we calculate the total number of skills we match to them, as well as the average value of these skills. Next, we estimate the following regression model:
@@ C.2 Salary versus skill value
where $w_{j}$ is the salary or midpoint of the salary range posted for job $j$, $\bar{v}_{j}$ is the average of the logarithm of the skill values across all skills that job $j$ requires, $\bar{r}_{j}$ the average of the skill density around these skills (i.e., the coherence of the skills the job requires) and $n_{j}$ the number of skills that job $j$ requires. $\tau_{y(j)}$ represents year fixed effects to control for general trends in the labor market. Finally, $\varepsilon_{j}$ is an error term.
Outcomes are reported in the table of Table C.2. The estimated effect effect of skill value on posted salaries is around 0.9 across all models, suggesting that a $10\%$ increase in skill value is associated with a $9\%$ higher salary.
@@
Robust standard errors in parentheses. p-value: ***: 0.01, **: 0.05, *: 0.1. The dependent variable is the value or midpoint of the salary range posted in a job ad, skill value is the average of the logarithm of the skill values across all skills the job requires, skill coherence is the average of the skill density round all skills the job requires, # skills is the logarithm of the number of skills the job requires.
C.3 Alternative skill matches and thresholds
@@
We test three alternative methods to match job requirements in job ads to skills. To do so, let $v_{t}^{j}$ denote the embedding vector of job requirement $t$ of job $j$. Our first method, used in the main text, calculates the cosine similarity between $v_{t}^{j}$ and the embedding vector of each skill’s (long) labels, and then selects the skill that yields the highest cosine similarity. Our second method uses the similarity of $v_{t}^{j}$ to the embedding of the most prominent tag (in terms of its usage on SO) in each skill. Our third method calculates the similarity of $v_{t}^{j}$ to the average embeddings of the tags defining a skill. Finally, our fourth method uses the similarity of $v_{t}^{j}$ to the embedding of the closest tag in each skill. Next, we repeat the analysis depicted in Fig. 2 with each of these methods. Fig. C.1 presents the results.
@@
Figure C.1: Alternative methods to match job requirements in job ads to skills. a,c,e,g. Prediction of skill requirements. $40\%$ of skill vector elements are masked and grouped into 10 equally sized bins based on the average relatedness of the masked skill to the (unmasked) skill requirements of a job. Plot displays the estimated probability, with its $95\%$ confidence interval, that skills in a bin are required in the job. b,d,f,h. Predictions of wage offers. Jobs are grouped into equally sized bins based on the average value of the skills they require. Vertical axis shows the estimated mean salary posted in the job ads, with their $95\%$ confidence intervals. a,b. Job requirements are matched to skills using the embedding vector of the skill’s label, namely the results shown in Fig 2. c,d. Job requirements are matched to skill using the embedding vector of the skill’s main (most frequent) tag. e,f. Job requirements are matched to skill using the average embedding vector across all of the skill’s tags. g,h. Job requirements are matched to skills using the closest embedding vector across all tags of a skill.
@@
In the main text, we only keep job requirements that can be matched to skills at a cosine similarity of $0.3$ or above. To explore the significance of this threshold, we repeat our analyses with two another thresholds: $0.2$ and $0.4$. The results are shown in Fig. C.2. Varying the threshold around the $0.3$ value of the main text does not meaningfully alter the outcomes of this analysis: density around masked skills strongly predicts which other skills a job demands and salaries posted in job ads rise with the average values of the skill we match to them. These alternative matching methods and thresholds show that outcomes are robust and do not depend on the LLM-generated skill labels.
@@
Figure C.2: Results of skill and salary predictions at varying matching thresholds. a,c,e. Probability that a masked skill is required in a job. Masked skills representing $40\%$ of skill vector elements are grouped into 10 equally sized bins based on the density around the masked skill of unmasked skill requirements in a job. Plot displays estimated probabilities in each bin. b,d,f Predictions of salaries. Jobs are grouped into equally sized bins based on the average value of their required skills. Vertical axis shows the estimated means in each bin of the advertised wage offers. a,b. Job requirements are matched to skills with a minimum cosine similarity of $0.3$, namely the results shown in Fig. 2. c,d. Job requirements are matched to skills with a minimum cosine similarity of $0.2$. e,f. Job requirements are matched to skills with a minimum cosine similarity of $0.4$. Vertical bars represent 95% confidence intervals.
@@ D.1 Alternative specifications for voting regression
In Fig. 4a of the main text, we study how the popularity of an answer relates to the skill expertise of the user that provided it. The figure shows that answers by users with more skill-specific expertise tend to receive more upvotes and are more likely to be the top answer to a question. Table D.1 provides the full set of regression results related to this analysis.
@@
Regression model of eq. (7). The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. skill expertise: logarithm of skill expertise, the total number of answers the users provided to questions related to the focal skill in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain either year or skill-year fixed effects.
Because for questions that receive only a single answer, this answer is automatically the top answer, we repeat the analysis in a sample that excludes any question that receives only one answer. Table D.2 shows that this leads to larger observed effects of skill expertise.
@@ D.2 Causal effects: Instrumental variable estimation
In the main text, we raise the issue of sample selection bias in the regression analysis of the number of upvotes an answer receives on SO. There are two concerns. First, users can observe all existing answers to a question and, arguably, users will take these answers into consideration before deciding whether or not to provide an answer themselves. As a consequence, the answers we observe on SO are no random sample of user-answer combinations. For instance, we expect that users who believe they have little to add to existing answers to a given question will tend to refrain from posting their own answers. This bias is addressed by only focusing on the first answer (in a temporal sense) that each question received.
@@
Second, even the first answer to a question will not be randomly sampled from a population of users and the answers they would have provided had they chosen to do so. Instead, answers will be provided by users who believe they have useful insights to the problem at hand. To address the sample selection bias this causes, we will exploit the fact that, due to timezone differences, most users are active on the SO platform at different, yet somewhat predictable, times. Therefore, we expect that who will provide the first answer to a question to some extent depends on the time of day the question was posted. This introduces some exogenous variation in the likelihood that a given user provides an answer to a given question. Moreover, as far as skill expertise is unequally distributed across timezones, this offers exogenous variation in the skill expertise associated with the first answer to a question.
@@
We proceed as follows. First, we label each minute of a day, henceforth day-minute, from 0 to 1399 in a manner that is independent of timezones. Next, for each user, we calculate the average minute of the day they post answers on SO. This offers a coarse indication of when a user is active on the platform. Finally, we calculate for each skill and each day-minute the amount of skill expertise that is expected to be available on SO. To do so, we sum the skill expertise of all users, weighted by the gap between the day-minute of a question and the average day-minute of each user. We call this variable, $M_{\theta,m}$, the minute-skill count of skill $\theta$ at day-minute $m$. Next, we use this variable as an instrument for the actual skill expertise of whichever user provides the first answer to the question at hand.
@@
Note that there are some potential concerns about this instrument. First, if there are geographic components to skill expertise of users and if these geographic components directly influence the quality of an answer, the instrument may be invalid. For instance, the timezone that includes Silicon Valley may have a disproportionate number of highly skilled programmers, such that any questions posed when most people in this timezone are active may attract high quality first answers. To correct for this, we always include day-minute fixed effects and in some specifications also add skill–year-combination fixed effects.
@@
Second, the instrument assumes that questions are answered within a day from the time they are posted on SO. In fact, we would expect that our instrument does not work for questions whose first answer does not arrive within the first 24 hours after the question was raised. We exploit this in placebo tests that focus on questions whose first answer is provided at least 24 hours after the posting of the question itself.
@@
where $u(a)$ is the user who posts answer $a$, $X_{\theta,u(a)}$ user $u(a)$’s skill-specific expertise in terms of the number of answers provided to questions involving skill $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers that will eventually be provided to question $q(a)$, $Q_{a}$ the set of these answers, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term. Compared to eq. (7) in the main text, we add day-minute fixed effects, $\mu_{(q_{a}})$, that control for any timezone specific confounders.
As before, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is the top answer to its question or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. If an answer is associated with multiple skills, we replicate the observation accordingly.
The results is presented in Table D.3. The effect of skill-specific expertise on answer popularity is significant and positive across all model specifications. Moreover, the estimated causal effects arrived at by instrumental variables estimation are substantially larger than the OLS estimates reported in the main text. This is in line with a downward sample-selection bias in the OLS results.
@@
Clustered (group_skill_qminute) standard-errors in parentheses
@@
Results of eq. (15) estimated with 2SLS. The dependent variable is either a binary variable that encodes whether or not an answer has received the most upvotes among all answers provided to a question, or the logarithm of the number of upvotes the answer received. skill expertise: logarithm of skill expertise, the total number of answers the user provided to questions related to the focal skill in the past two calendar years. # answers: logarithm of the total number of answers provided to the focal answer’s question. # votes all answers: logarithm of the total number of votes received by these answers. Regression models contain day-minute fixed effects (QA-abs-minute) for the timing of the question and year or -year fixed effects. To overcome sample-selection biases, we instrument skill expertise with $M_{\theta,m}$, the estimated amount of available skill expertise at the moment the question was posted, calculated as the weighted sum of skill expertise across all users, where weights reflect the temporal distance between the day-minute of a user and of the question’s posting.
Next, we test this strategy using our placebo tests. To do so, we split the sample of questions into questions that receive a first answer within the first 24 hours after they were posted and those that receive their first answer later than that. Results are shown in Table D.4. For answers provided within 24 hours, skill expertise has a strong and statistically significant positive effect on answer popularity, regardless of whether we look at the probability of being the top answer or the number of upvotes an answer receives. In contrast, for the placebo test, comprising of answers provided more than 24 hours after the question was posted, the effect of skill expertise becomes insignificant in three out of four models. In the model without skill fixed effects, the effect on the number of votes an answer receives turns significantly negative. This suggests that the skill fixed effects control for some relevant confounders. Therefore, our preferred estimates are derived from models that include these fixed effects. These models suggest a large and positive effect of skill expertise on the popularity of answers, corroborating our analysis in the main text.
@@
Clustered (group_skill_qminute) standard-errors in parentheses
@@
D.3 Regression of skill acquisition on skill value
To study how skill values effect which new skills users learn, we run the regression described in eq. (9) of the main text. We compare the association of two variables, skill value and skill density, with the likelihood that users acquire a certain new skoill. We repeat this analysis once using all answers to construct expertise and skill acquisition variables, and once using only answers to Python related questions. The results are shown in Table D.5.
Table D.5: Skill acquisition regressions
Dependent Variable: skill acquisition (binary, all programming languages) skill acquisition (binary, Python)